### v1

# Importance Guided Data Augmentation for Neural-Based Code Understanding

Abstract
Pre-trained code models lead the era of code intelligence. Many models have been designed with impressive performance recently. However, one important problem, data augmentation for code data that automatically helps developers prepare training data lacks study in the field of code learning. In this paper, we introduce a general data augmentation framework, GenCode, to enhance the training of code understanding models. GenCode follows a generation-and-selection paradigm to prepare useful training codes. Specifically, it uses code transformation techniques to generate new code candidates first and then selects important ones as the training data by importance metrics. To evaluate the effectiveness of GenCode with a general importance metric – loss value, we conduct experiments on four code understanding tasks (e.g., code clone detection) and three pre-trained code models (e.g., CodeT5). Compared to the state-of-the-art (SOTA) code augmentation method, MixCode, GenCode produces code models with 2.92% higher accuracy and 4.90% robustness on average.
1 Introduction
Deep learning (DL) for code has been one of the successful DL applications over the years. The methodology of applying deep learning to solve software engineering (SE) tasks, such as software vulnerability detection Zhou et al. (2019), is commonly known as code learning. From the model level, many DL models specifically designed for code learning have been proposed, such as CodeBERT Feng et al. (2020), GraphCodeBERT Guo et al. (2020), and CodeT5 Wang et al. (2021). From the business perspective, multiple related products such as Copilot  [^1] https://github.com/features/copilot and Tabnine  [^2] https://www.tabnine.com have been widely used to help our daily development. As a result, in different downstream learning-based code understanding tasks, e.g., code clone detection  Lu et al. (2021) and program classification Puri et al. (2021), DL already achieved significantly better results than traditional SE techniques.
However, most of the existing research focuses on designing new code representation techniques or code model architectures to further improve the effectiveness of DL for code, one important problem – how to prepare high-quality training data gained limited attention even though recent research demonstrated it highly affects the performance of code models Sun et al. (2022). The challenge behind this is preparing training data for code models is heavy work. Although today developers can easily collect raw code data from the open-source community such as GitHub [^3] https://github.com, data labeling is still expensive, especially for code-related tasks. It needs human effort with professional domain knowledge.
Techniques can be directly used for code augmentation, such as code refactoring methods and text transformation methods. Unfortunately, existing studies Bielik and Vechev (2020); Yu et al. (2022) stated that code refactoring methods have limited advantages in boosting the training of code models. Besides,  Dong et al. (2023b) revealed that almost all the data augmentation methods cannot produce code models with good robustness. Moreover, a recent work, MixCode Dong et al. (2023a), that follows the spirit of Mixup Zhang et al. (2018) to do code augmentation achieved promising results, but it can be only used for classification tasks. Therefore, there are no effective and reliable code augmentation methods that can be generalized to different tasks.
To bridge this gap, inspired by the great success of the typical two-stage Automated Data Augmentation (AutoDA) framework in the field of computer vision Cubuk et al. (2020); Yang et al. (2023b), we propose GenCode, the first generation-and-selection-based data augmentation framework for code understanding. Specifically, at each training epoch, GenCode first generates training candidates by using different code transformation methods. Here, we consider different types of code transformation methods such as semantic-preserving code refactoring methods and syntax-breaking text transformation methods. After that, GenCode selects valuable training data from the candidates with the guidance of importance metrics. The important metric used in GenCode is simply defined as the loss value of code data produced by the code model. Finally, the selected samples are used for the model training. We conducted comprehensive experiments to study how GenCode can improve the accuracy and robustness of code models. The experimental results demonstrated that, compared to the SOTA method, MixCode, GenCode can produce code models with a 2.92% higher accuracy and 4.90% better robustness on average.
To summarize, the main contributions of this paper are:
  • We propose the first generation-and-selection-based framework, GenCode, which automatically enhances the training data and identifies the most important data from the augmented set for boosting the code understanding.
  • We empirically demonstrate that GenCode can produce more accurate and robust code understanding models compared to existing data augmentation methods and conduct an ablation study to explore how different importance scores affect the effectiveness of GenCode.
2 Background and Related Work
2.1 Deep Learning for Code
Since source code can be seen as natural language or structural graph Allamanis et al. (2018), researchers employ the techniques in Natural Language Processing (NLP) and Graph Neural Networks (GNNs) to learn information from code and solve downstream tasks, such as automated bug detection Dinella et al. (2020), automated program classification Puri et al. (2021), and code clone detection Wang et al. (2020). The key problem in deep learning for code is how to represent raw code to machine-readable inputs. Researchers designed different methods for code representation that can be roughly divided into two groups, sequence-based representation and graph-based representation.
Sequence-based representation follows the same procedure of text data processing in NLP to convert raw code data into a sequence of code tokens. For example, given a code snippet, “def func(x, y)”, sequence-based representation transforms it to a token list as “[’def’, ’func’, ’(’, ’x’, ’y’, ’)’]”. After that, each token in the token list will be mapped to a pre-defined index value, e.g., [1, 25, 34, 5100, 640, 33]. Finally, these index values will be fed into the code-embedding models for the learning.
Graph-based representation learns the structural information of code. Here, the structural information can be abstract syntax tree (AST), data flow information, and other types of flow information Dinella et al. (2020); Yasunaga and Liang (2020). For example, GraphCodeBERT combines the sequence-based representation and data flow information of code as a whole to train code models with better performance.
In this work, we focus on both sequence-based and graph-based representation and study the powerful pre-trained code models CodeBERT, GraphCodeBERT, and CodeT5 that achieved impressive results on code understanding tasks.
2.2 Data Augmentation for Code
Data augmentation is a well-studied technique in the machine learning field to increase the size and diversity of training data by only modifying the existing labeled data, thus, without further human effort. For instance, in the NLP field, researchers tried many ways to augment the text data, e.g., replacing a token with a synonym Feng et al. (2021).
Recently, data augmentation gained attention from software engineering researchers. Various data augmentation methods designed for source code have been proposed Gao et al. (2020); Yefet et al. (2020); Allamanis et al. (2021); Pour et al. (2021); Bui et al. (2021); Wang et al. (2022); Park et al. (2023); Yang et al. (2023a), with the most general framework involving the use of code refactoring, e.g., local variable renaming, duplication, and dead store, to create more code data without breaking the semantics of the original codes. Concretely, code refactoring rewrites the syntactic structure of source code while keeping the semantic information Lacerda et al. (2020).
Existing works Bielik and Vechev (2020); Yu et al. (2022); de Paula Rodrigues et al. (2023); Dong et al. (2023b, 2024) conducted empirical studies and found code refactoring methods have limited advantages in boosting the training of code models. For instance, a recent empirical study Dong et al. (2024) on code understanding revealed that the utilization of code refactoring methods only achieved success rates of 66.67% and 41.67% in improving accuracy and robustness, respectively, compared to code model training without data augmentation. Besides, the SOTA code data augmentation method, MixCode which linearly interpolates pairs of source codes, has been proven more effective than the current general framework code refactoring. However, as a variant of Mixup specifically designed for code data, MixCode is limited to applications in code classification tasks, as it necessitates the label with one-hot encoding.
Compared to existing works, we propose a simple yet effective code augmentation framework that is not only unaffected by the dataset and model constraints but also applicable to both classification and non-classification tasks.
3 Definition
We first define the problem of data augmentation for code models.
Definition 3.1 (Program Code [^4] For simplicity, we use code in the paper. ).
A code is defined as a 5-tuple $\mathcal{P}=\left(content,se,sy,Ip,Op\right)$: $content$ contains all the text in the code $\mathcal{P}$; $se$ and $sy$ are the semantic and syntax of the code $\mathcal{P}$; and $Ip$ and $Op$ are the input and output of the code $\mathcal{P}$.
Definition 3.2 (Code Model).
A code model is defined as a 3-tuple $M=\left(parameter,\mathcal{P},Y\right)$, where $parameter$ contains all the weights in code model $M$, $\mathcal{P}$ is the input code, and $Y$ is the label of $\mathcal{P}$ given a specific downstream code task.
Code models can be any machine learning model such as a simple regression model or a pre-trained model, e.g., CodeT5 Wang et al. (2021).
Definition 3.3 (Code Augmentation).
Given a code $P$, a code transformation method $\mathcal{T}:\mathcal{X}\rightarrow\mathcal{X^{\prime}}$, code augmentation is to transform $\mathcal{P}$ to $\mathcal{P^{\prime}}$ that $\mathcal{P^{\prime}}.content=\mathcal{T}\left(\mathcal{P}.content\right)$.
Here, code transformation methods can be any method that can change the content of program statements such as text data transformation methods (e.g., synonym replacement and code refactoring methods (e.g., dead if statement adding).
Definition 3.4 (Semantic-preserving/breaking Code Data Augmentation).
Given a code model $M$, a code $\mathcal{P}$ and its augmented version $\mathcal{P^{\prime}}$, semantic-preserving (breaking) code augmentation indicates $\mathcal{P}.se=\mathcal{P^{\prime}}.se$ $M(\mathcal{P})=M(\mathcal{P^{\prime}})$ ($\mathcal{P}.se\neq\mathcal{P^{\prime}}.se$ $M(\mathcal{P})\neq M(\mathcal{P^{\prime}})$).
Definition 3.5 (Syntax-preserving/breaking Code Augmentation).
Given a code $\mathcal{P}$ and its augmented version $\mathcal{P^{\prime}}$, syntax-preserving (breaking) code augmentation indicates $\mathcal{P}.sy=\mathcal{P^{\prime}}.sy$ ($\mathcal{P}.sy\neq\mathcal{P^{\prime}}.sy$).
4 Assumption and Preliminary Study
GenCode selects the most valuable data for the code model training. This approach is grounded on the assumption that the data that code models have less confidence in is more useful for model training. To check whether this assumption is valid or not, we conducted a preliminary study to explore if there is a correlation between the accuracy of the trained code model and the confidence of the model on the training data. Specifically, the study is conducted on the Google Code Jam (GCJ) Yang et al. (2022) dataset and CodeBERT pre-trained code model. The detailed steps are as follows:
Step 1. We initialize the code model and use code augmentation methods to generate additional data as the training candidate set. The code augmentation methods will be introduced in Section 5.2.
Step 2. We randomly select a number of data from the training candidate set. Here the selected number is the same as the size of the original training data. We repeat this selection 100 times and collect 100 groups of training data. After that, for each group of training data, we compute and record their average loss values.
Step 3. We train the code model using each group of training data respectively for a few epochs and thus save 100 trained code models. We then test each code model using the original test data and save 100 accuracies.
Step 4. We use the Pearson Correlation Coefficient (PCC) Schober et al. (2018) metric to check the correlation between the loss values and the accuracies.
Table 1: Correlation between loss values and code model accuracy.
Coefficient P-value
Initialized (epoch 0) 0.6075 7.9687e-12
Middle (epoch 10) 0.2440 0.0125
Well-trained (epoch 20) 0.1657 0.0926
Table 1 summarizes the results of our preliminary study. Here, we considered three types of code models, the initialed model, the model that has been trained for a few epochs but not converged, and the well-trained code model, to check if our assumption can stand in different scenarios. The results indicate that there is a positive correlation between the loss values of data and the accuracy of code models trained by using these data. From the initialized model to the well-trained model, this correlation becomes weaker but still exists. Therefore, we can propose an effective and reliable data augmentation framework to boost the training of code models based on this phenomenon.
Finding: There is a positive correlation between the loss values of code data and the accuracy of code models trained by using these data. When the code model is randomly initialized, this correlation is significant (with a P-value less than 0.01).
5 Methodology
5.1 Overview
The goal of GenCode is to produce code models with better performance on test data by training models on carefully generated data. Given a training set of program $\mathcal{P}_{train}$, its transformed version $\mathcal{P^{\prime}}_{train}$, and a set of test programs $\mathcal{P}_{test}$, after the training process $\mathsf{train}\left(\cdot\right)$ , $\varrho(M^{\prime}(\mathcal{P}_{test}))>\varrho(M(\mathcal{P}_{test}))$, where $M.parameter=\mathsf{train}(M,\mathcal{P}_{train})$ $M^{\prime}.parameter=\mathsf{train}(M^{\prime},\mathcal{P^{\prime}}_{train})$, and $\varrho\left(\cdot\right)$ is the performance measurement.
GenCode follows the generation-and-selection workflow and utilizes different types of code transformation methods to generate diverse code data first, then selects code data that are most useful for training as the training data. Algorithm 1 shows how GenCode works. In each training epoch, GenCode automatically generates more code data (Line 7 to Line 9, Definition 3.3) and then selects the most important ones (Line 10 to Line 13) to train the code model (Line 14). Here, the data importance is measured by the loss values produced by the code model which is the most common indicator to check how the model performs on the data. Overall, GenCode contains two main steps, utilizing code data augmentation methods to compose the search space and importance-guided data selection.
Algorithm 1 GenCode
1:$M$: initialized DNN model
2:$\mathcal{P},Y$: original code data and labels
3:$\mathcal{T}=\left\{t\right\}:$ a set of code augmentation methods
4:$M$: trained model
5:for $run\leftarrow\{0,\ldots,\#epochs\}$ do
6:     $\mathcal{P^{\prime}},Y^{\prime}\ \leftarrow\phi,\phi$triangleright initialization
7:     for $t\in\mathcal{T}$ do
8:         $\mathcal{P^{\prime}}\leftarrow\mathcal{P^{\prime}}\cup t\left(\mathcal{P}\right)$
9:         $Y^{\prime}\leftarrow Y^{\prime}\cup Y$triangleright code augmentation
10:     $I\leftarrow\mathsf{Importance}\left(M,\mathcal{P^{\prime}}\right)$triangleright computing importance
11:     $Index\leftarrow\mathsf{argSort}\left(I\right)\left[:||\mathcal{P}||\right]$triangleright sort by importance
12:     $X_{train}\leftarrow\mathcal{P^{\prime}}\left[Indext\right]$
13:     $Y_{train}\leftarrow Y^{\prime}\left[Indext\right]$
14:     $M\leftarrow\mathsf{Train}\left(M,X_{train},Y_{train}\right)$
15:return $M$
5.2 Search Space
Even though the code has strict syntax constraints, recent work Dong et al. (2023b) found that these data augmentation methods that slightly break the syntax of the source code e.g., Random Swap, can still be useful for code model training. To enhance the diversity of source code within the feature space, therefore, GenCode considers two types of code augmentation methods, semantic/syntax-preserving methods, and syntax-breaking methods (see Definition 3.4 and Definition 3.5).
GenCode adapts code refactoring methods as semantic/syntax-preserving methods. Code refactoring methods follow the constraints of programming languages to edit the code while keeping the main functionality of the code. On the other hand, text augmentation methods are used as syntax-breaking methods in GenCode which are widely used in the NLP field. For example, text augmentation methods can randomly delete some tokens in the code. However, even though the syntax has been broken, the code model can still successfully predict those codes since it reads the code as a sequence. In total, as shown in Table 2, GenCode currently supports 18 code refactoring methods and 5 text augmentation methods. Besides, Figure 1 provides an example of the training candidate that is to generate the search space. In the syntax-preserving, we present a code refactoring method named API Renaming, which involves modifying the function name from ”main()” to ”even()”. In the syntax-breaking, we demonstrate another instance of API Renaming, where two different statements within a program are randomly selected and their positions swapped. As depicted in this example, the third and fourth lines have been interchanged. Appendix B detailed introduces the specifics of these atomic data augmentation operators.
Table 2: Atomic data augmentation operators adopted in GenCode.
Syntax-preserving Syntax-breaking
API renaming
Filed enhancement
Synonym replacement Random insertion Random swap Random deletion Back-translation
Arguments adding
For loop enhancement
Arguments renaming
If enhancement
Dead for adding
Local variable adding
Dead if adding
Local variable renaming
Dead if-else adding
Method name renaming
Dead switch adding
Plus zero
Dead while adding
Print adding
Duplication Return optimal
Figure 1: An example of atomic data augmentation operators.
Table 3: Effectiveness of data augmentation methods w.r.t. test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on original test data and ASR $\downarrow$ (%) on test data. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
GCJ Refactory Java250 BigCloneBench
Model DA method Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$
No Aug 90.96 ± 0.07 33.31 56.13 96.21 ± 0.13 36.31 42.31 96.36 ± 0.09 40.55 53.37 96.91 ± 0.12 9.74 24.61
Refactor 91.74 ± 0.11 35.56 52.84 95.95 ± 0.19 32.64 40.89 96.44 ± 0.12 38.35 52.56 96.97 ± 0.21 8.55 18.29
SR 70.66 ± 0.11 35.46 52.71 96.71 ± 0.11 39.25 53.91 96.21 ± 0.02 42.01 53.79 97.02 ± 0.15 12.52 26.11
RI 59.91 ± 0.21 40.78 65.76 96.43 ± 0.18 38.45 45.57 96.15 ± 0.06 43.15 55.07 97.38 ± 0.23 8.31 30.08
RS 93.24 ± 0.13 28.26 52.44 96.73 ± 0.15 38.63 45.98 96.49 ± 0.04 39.36 50.79 96.92 ± 0.12 13.41 30.13
RD 76.98 ± 0.09 35.65 55.46 96.46 ± 0.18 30.23 41.32 96.61 ± 0.12 38.31 52.42 97.11 ± 0.19 9.58 30.15
BT 82.74 ± 0.15 40.38 55.95 94.31 ± 0.38 62.98 69.87 96.06 ± 0.06 42.21 53.32 97.13 ± 0.17 11.29 26.32
MixCode 92.21 ± 0.25 29.28 52.53 96.76 ± 0.31 31.04 41.01 96.62 ± 0.23 36.14 50.67 - - -
CodeBERT GenCode 96.23 ± 0.18 22.33 46.62 99.32 ± 0.14 23.75 31.25 98.89 ± 0.09 29.89 40.65 98.93 ± 0.11 7.12 15.83
No Aug 93.96 ± 0.15 31.23 56.02 96.84 ± 0.16 22.97 30.26 96.41 ± 0.11 23.16 42.66 96.88 ± 0.09 4.19 12.22
Refactor 91.72 ± 0.21 30.31 56.58 95.49 ± 0.21 24.26 32.57 96.53 ± 0.15 23.13 42.42 97.11 ± 0.23 6.05 6.17
SR 89.46 ± 0.13 40.02 66.68 95.57 ± 0.13 28.39 47.15 96.46 ± 0.14 22.92 42.18 97.14 ± 0.16 6.13 10.16
RI 80.47 ± 0.21 43.51 65.75 94.51 ± 0.25 25.08 32.87 96.52 ± 0.09 22.37 42.33 97.35 ± 0.17 2.23 6.28
RS 94.76 ± 0.15 30.18 54.75 97.88 ± 0.15 21.85 29.67 96.44 ± 0.14 22.68 41.98 97.54 ± 0.12 6.03 11.94
RD 91.71 ± 0.26 31.73 60.92 96.53 ± 0.19 20.35 26.98 96.56 ± 0.21 22.39 41.84 97.16 ± 0.21 6.18 8.27
BT 80.47 ± 0.05 41.65 62.99 95.99 ± 0.27 35.56 46.23 96.39 ± 0.13 22.91 42.32 97.11 ± 0.16 8.18 14.12
MixCode 94.19 ± 0.25 30.26 54.49 97.29 ± 0.15 14.89 26.53 96.89 ± 0.19 20.03 42.01 - - -
GraphCodeBERT GenCode 97.93 ± 0.14 26.43 48.87 99.45 ± 0.17 15.78 21.92 99.16 ± 0.13 18.76 33.81 99.23 ± 0.08 1.86 4.62
No Aug 95.02 ± 0.07 30.78 55.67 97.22 ± 0.17 22.45 29.69 96.98 ± 0.08 31.76 49.82 97.38 ± 0.13 10.32 22.97
Refactor 92.21 ± 0.14 30.12 55.93 96.43 ± 0.22 28.76 35.45 96.92 ± 0.11 30.66 47.12 97.24 ± 0.12 9.32 22.05
SR 90.14 ± 0.21 39.21 51.84 94.67 ± 0.14 28.24 36.26 96.53 ± 0.05 28.65 46.01 97.31 ± 0.08 15.87 25.98
RI 78.77 ± 0.05 41.02 63.46 95.02 ± 0.15 24.78 33.61 96.11 ± 0.09 27.14 46.76 97.46 ± 0.13 10.43 29.76
RS 95.66 ± 0.14 29.35 51.74 97.97 ± 0.29 21.32 27.76 96.34 ± 0.11 28.45 48.45 97.03 ± 0.15 14.33 30.43
RD 92.01 ± 0.11 31.04 56.85 97.52 ± 0.26 19.93 24.87 97.21 ± 0.07 26.09 45.24 97.43 ± 0.07 9.98 28.87
BT 81.45 ± 0.13 39.65 54.24 95.38 ± 0.11 33.62 43.67 96.45 ± 0.14 29.47 45.38 97.65 ± 0.14 10.43 25.34
MixCode 94.32 ± 0.16 29.82 42.56 97.32 ± 0.13 12.49 18.32 96.91 ± 0.11 28.73 48.87 - - -
CodeT5 GenCode 98.85 ± 0.11 23.57 44.76 99.64 ± 0.06 12.13 19.56 99.38 ± 0.11 21.63 39.76 99.34 ± 0.12 7.39 12.15
5.3 Importance-Guided Data Selection
After constructing the training candidate (as introduced in Section 5.2), the next step is to choose the most valuable code data that can benefit the code model training. Based on our Finding drawn from the preliminary study in Section 4, GenCode chooses the simplest way to measure the importance of code data to the model – the loss value. That means, GenCode takes all the transformed code as input to the code model and records all the loss values first, then ranks all the code in descending order based on the loss values and selects the first $K$ code as the training data. Following the common setting in the data augmentation problem Dong et al. (2023a), $K$ is set as the number of original training data. In GenCode, we choose loss value as the indicator since it can be generalized to any code model and task.
6 Evaluation
We evaluate both the accuracy (Section 6.2) and robustness (Section 6.3) of trained code models. Besides, we check how does importance score affects the effectiveness of GenCode (Section 6.4).
6.1 Datasets and Models
Datasets and Models We evaluate GenCode on four code understanding tasks, bug detection Hu et al. (2019), authorship attribution Yang et al. (2022), code clone detection Yang et al. (2022) (Note that the SOTA method, MixCode, is not suitable for this task since its input is code pairs.), and problem classification Puri et al. (2021). For each code task, we fine-tune three types of models based on pre-trained code models, CodeBERT, GraphCodebERT, and CodeT5. For the model training, we follow existing works Hu et al. (2019); Wang et al. (2021); Puri et al. (2021); Yang et al. (2022); Dong et al. (2023a), to set all configurations. Specifically, we set the training epoch as 50 for all tasks and models. Regarding the optimizer, we utilize Adam Kingma and Ba (2014) with a learning rate of $10^{-3}$ for all the models. Besides, we use early stopping with a patience of 20 to mitigate overfitting. To reduce the influence of randomness, we train each model five times and report the average results along with the standard deviation. For MixCode, we follow the recommendation of the original paper and set the Mixup ratio as $\alpha=0.1$. For the setting of adversarial attacks, we directly use the default configurations provided by the original papers  Yang et al. (2022) in our work. More details are explained in Appendix A and C.
Evaluation Metrics. In the evaluation, we assess the performance of trained models from two perspectives, namely, clean accuracy and model robustness. For the robustness evaluation, we check the attack success rate (ASR) metric of two well-established adversarial attack methods, namely, ALERT Yang et al. (2022) and MHM Zhang et al. (2020) on code models. We do not consider the natural robustness since the existing natural robustness datasets can not fully support our studied tasks, i.e., no natural robustness datasets support authorship attribution, bug detection, and code clone detection.
Baseline Methods. We consider model training without data augmentation (No Aug), Refactor, a family of text-oriented data augmentation methods, and MixCode as baseline methods. Refactor is the widely used general data augmentation framework that has been specifically designed for code data, while MixCode is the more recent code data augmentation method with high performance in enhancing code models Dong et al. (2023a). These baselines allow us to determine if there are real improvements in the accuracy and robustness of the code model when using GenCode.
6.2 Accuracy Evaluation
Overall accuracy. Column Accuracy in Table 3 summarizes the final test accuracy of code models trained by using different code data augmentation methods including GenCode. Surprisingly, from the results, we can see that most of the existing code augmentation methods except RS and MixCode negatively affect the performance of code models. This finding is similar to the conclusion from recent code augmentation studies Bielik and Vechev (2020); Yu et al. (2022); Dong et al. (2023b) which states code augmentation has limited advantage in boosting the training of code models and also indicates that data augmentation for code is a challenging problem.
Then, compared to the existing data augmentation methods, the results demonstrate that GenCode consistently enhances the training and achieves the best results across all the considered datasets and models. Concretely, on average, GenCode generates models with 4.52%, 3.05%, 2.56%, and 2.11% higher accuracy than No Aug in GCJ, Refactory, Java250, and BigCloneBench, respectively. Besides, compared to the SOTA code augmentation method MixCodde, GenCode can achieve 4.1%, 2.34%, and 2.33% better results in GCJ, Refactory, and Java250 datasets, respectively. In conclusion, GenCode is effective in training accurate code understanding models.
(a) Authorship attribution
(b) Bug detection
(c) Problem classification
(d) Clone detection
Figure 2: Convergence speed of CodeBERT using different code data augmentation methods in each task.
Case study 1: Does the training epoch affect the accuracy of GenCode? Figure 2 depicts the average test accuracy at each training epoch of CodeBERT using different data augmentation methods. The results demonstrated that GenCode consistently outperforms other baseline methods during the whole training process. We can conclude that the choice of the number of training epochs will not affect the advantages of our method over other methods.
(a) Refactor
(b) SR
(c) RI
(d) RS
(e) RD
(f) BT
(g) MixCode
(h) GenCode
Figure 3: Visualization of code embeddings after dimension reduction using Principal Component Analysis (PCA). Model: CodeBERT, dataset: Refactory, task: Bug detection.
Table 4: A case study of mean values of maximum probabilities. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
Model DA method GCJ Refactory Java250 BigCloneBench
No Aug 0.0381945 0.9368535 0.9554731 0.9651272
Refactor 0.0391864 0.9643155 0.9620814 0.9654681
SR 0.0380562 0.9527686 0.9341456 0.9583812
RI 0.0371215 0.9189242 0.9327171 0.9704512
RS 0.0392987 0.9561591 0.9581767 0.9410937
RD 0.0392987 0.9723022 0.9634647 0.9631375
BT 0.0379581 0.9566747 0.9417861 0.9667154
MixCode 0.0390714 0.9687351 0.9685594 -
CodeBERT GenCode 0.0413779 0.9808122 0.9714346 0.9786421
Table 5: Effectiveness of data augmentation methods w.r.t. test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on original test data and ASR $\downarrow$ (%) on test data. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
GCJ Refactory Java250 BigCloneBench
Model DA method Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$ Accuracy normal-$\uparrow$ MHM normal-$\downarrow$ ALERT normal-$\downarrow$
GenCode (Random) 94.74 ± 0.21 25.21 50.32 98.82 ± 0.15 27.65 36.65 98.24 ± 0.11 32.52 44.87 98.27 ± 0.19 8.01 18.12
GenCode (Min) 93.34 ± 0.32 35.66 66.67 96.13 ± 0.21 65.51 69.76 96.43 ± 0.17 46.76 53.86 97.23 ± 0.14 20.65 41.32
CodeBERT GenCode 96.23 ± 0.18 22.33 42.62 99.32 ± 0.24 22.75 29.25 98.89 ± 0.09 29.89 40.65 98.93 ± 0.15 7.12 15.83
GenCode (Random) 95.26 ± 0.11 29.97 50.63 98.93 ± 0.18 17.56 24.29 98.68 ± 0.22 20.65 38.86 98.67 ± 0.13 2.04 5.97
GenCode (Min) 93.53 ± 0.17 40.55 68.23 96.31 ± 0.24 38.73 46.98 96.72 ± 0.12 26.87 46.67 97.22 ± 0.21 11.45 21.23
GraphCodeBERT GenCode 97.93 ± 0.14 26.43 39.87 99.45 ± 0.17 15.78 20.92 99.16 ± 0.13 18.76 33.81 99.23 ± 0.09 1.86 4.62
GenCode (Random) 96.24 ± 0.21 25.03 47.61 98.99 ± 0.21 18.46 23.67 98.93 ± 0.16 25.11 42.34 98.83 ± 0.08 8.67 17.42
GenCode (Min) 94.11 ± 0.31 38.65 59.89 98.29 ± 0.12 35.67 48.59 97.03 ± 0.23 32.61 49.27 97.01 ± 0.15 18.36 25.54
CodeT5 GenCode 98.85 ± 0.16 23.57 38.76 99.64 ± 0.37 12.13 18.81 99.38 ± 0.11 21.63 39.76 99.34 ± 0.12 7.39 12.15
Case study 2: What factors contribute to the improved accuracy observed in code models when utilizing GenCode? Furthermore, we explore why models trained by GenCode have higher accuracy than models trained by using other code data augmentation models. Specifically, 1) first, we extract the code embeddings produced by each trained model using all the test data as code features. 2) Then, since the extracted features are high dimensions and hard to analyze, we reduce the dimensions of code data features using the Principal Component Analysis (PCA) Maćkiewicz and Ratajczak (1993) algorithm. 3) After that, we visualize the reduced 2-dimensional vectors. Normally, the features extracted by a well-trained model are more clearly distributed based on the data labels. Figure 3 depicts the visualized features extracted from CodeBERT using the Refactory dataset. We can see that the code embeddings trained by using GenCode are more clearly distributed than embedding extracted by other models, i.e., the buggy code and correct code are greatly different in the figure. These results indicate that GenCode can train good code embeddings and, therefore, produce better accurate code models for code understanding tasks.
6.3 Robustness Evaluation
Overall robustness. Column MHM and ALERT in Table 3 summarize the attack success rate of code models. First, we can see that similar to the finding in Section 6.2, not all data augmentation methods can improve the robustness of code models. Specifically, in the baseline methods, only RS and MixCode have a positive impact on the robustness of code models. Considering our method, we can see that GenCode consistently improves the robustness of code models and achieves the best results in most cases. On average, compared to No Aug, models trained by GenCode have an 8.42% lower ASR (higher robustness). This result breaks the previous finding Bielik and Vechev (2020) that states simply adding additional generated code to the training dataset to train code models has limited benefits in improving their robustness. On the other hand, compared to the SOTA data augmentation method MixCode, GenCode can train code models with 4.90% higher robustness on average. In conclusion, GenCode can produce more robust code understanding models than other methods.
Case study 3: What factors contribute to the improved robustness observed in code models when utilizing GenCode? To explore why GenCode produces a more robust code model, we check the confidence of trained models using different code data augmentation methods on the original test data. Normally, if the DNN model is more confident in the test data, it is more difficult to attack this model using this data Wu et al. (2018). To do so, for each dataset and code model, we collect the maximum probability of the corrected classified test data produced by the code model first. Then, we calculate the mean values of these maximum probability scores. The results of the mean values of maximum probabilities are shown in Table 4. We can see that code models trained by GenCode are more confident in the test data than models trained using other code data augmentation methods.
6.4 Ablation Study
Even though we have preliminarily shown that selecting the code samples with higher loss values is a valid way to train good models in Section 4. It is still necessary to study to what extent the advantage can be brought from this selection strategy. Therefore, in this part, we conduct an ablation study to compare GenCode with its two variants, GenCode with data selection using minimum loss values and GenCode with random selection, GenCode (Min) and GenCode (Random), respectively.
Table 5 summarizes the accuracy and robustness of trained code models using different selection strategies. We can see GenCode consistently outperforms the two variants in both testing criteria. Selecting code samples with minimum loss values is the worst way for GenCode. Interestingly, we found that GenCode (Random) performs better than all baseline methods (i.e., No Aug, Refactor, and MixCode) in terms of both accuracy and robustness improvements based on all results. This indicates simply combining all syntax-preserving methods including code refactoring methods and syntax-breaking including text-oriented methods is already a good code data augmentation framework.
7 Conclusion
We proposed the first generation-and-selection-based framework, GenCode, for boosting code understanding models. The basic assumption behind GenCode is that the less confident data contributes more to the model training. Based on this assumption, GenCode utilizes atomic data augmentation operators to generate candidate training data first and selects data with maximum loss values to train the code model. Our evaluation demonstrated that GenCode can produce more accurate and robust models than MixCode with 2.92% accuracy and 4.90% robustness improvement, respectively. Besides, we analyzed how the importance score affects the effectiveness of GenCode.
In the future, we plan to 1) explore the combination of GenCode with other confidence scores, such as uncertainty metrics and active learning methods; 2) study the effectiveness of GenCode on fine-tuning open-source large language models, e.g., Llama.
Appendix A Datasets and Models
We evaluate GenCode on four code understanding tasks and three pre-trained code models as listed in Table 6. For all datasets, we follow the settings provided by the official projects for splitting the data into training, validation, and test sets.
  • Bug detection in source code understanding involves determining whether a piece of code contains bugs. This task can be thought of as a binary classification problem, where the target is to classify code as either bug-free or containing bugs. In our study, we utilize the Refactory dataset Hu et al. (2019), which is specifically designed for bug repair tasks. Refactory comprises 2,242 correct and 1,783 buggy Python programs. These programs were written by real-world undergraduate students to finish five different programming assignments.
  • Authorship attribution is the identification of the writer of a specific code fragment by inferring the unique characteristics of programmers from their published source code. This task serves important purposes, such as properly crediting a programmer for their contributions and aiding in the detection of code plagiarism. In our experiment, we employ a dataset sourced from Google Code Jam (GCJ) that is provided by the work Yang et al. (2022) for the authorship attribution task.
  • Clone detection is the task that focuses on determining whether two pieces of code are semantically identical. This task is instrumental in preventing the propagation of bugs and simplifying the process of software maintenance. To conduct our experiment, we adopt the widely acknowledged clone detection benchmark dataset BigCloneBench Svajlenko et al. (2014). This dataset offers an extensive and diverse set of code examples that are written by Java. Note that MixCode is not suitable for this task since its input is code pairs.
  • Problem classification is a typical task in source code understanding. It involves classifying the target functions of source code. Specifically, in a series of problems including detailed descriptions as well as their corresponding candidate source code, the trained model can correctly associate the source code with the specific problem that the code is trying to address. In our study, we leverage the recently released dataset Java250 Puri et al. (2021). Java250 is curated explicitly for Java program classification, containing a diverse set of 250 classification problems. Each problem within this dataset is paired with 300 different Java programs, each for solving this problem.
For the models, we study the powerful pre-trained code models.
  • CodeBERT Feng et al. (2020) is a bimodal model that is trained by leveraging data from various programming languages like C, C++, and Java, along with natural languages. Inspired by the principles of BERT Devlin et al. (2018), CodeBERT treats programs as sequences during pre-training.
  • GraphCodeBERT Guo et al. (2020) enhances the training data with data-flow information to incorporate the semantic-level structure of programs. This addition enriches the training process, enabling the model to generate more precise and context-aware code representations.
  • CodeT5 Wang et al. (2021) is a pre-trained encoder-decoder Transformer model that is designed to effectively capture the code semantics conveyed through developer-assigned identifiers.
Table 6: Details of tasks, datasets, and DNN models. #Training and #Test are the number of data for training and testing, respectively.
Dataset Language Task #Training #Test Model
Google Code Jam (GCJ) Python Authorship attribution 528 132 CodeBERT GraphCodeBERT CodeT5
Refactory Python Bug detection 3,380 423
Java250 Java Problem classification 48,000 15,000 CodeBERT GraphCodeBERT CodeT5
BigCloneBench Java Clone detection 90,102 4,000
Appendix B Atomic data augmentation operators
Code Refactoring (Refactor). Code refactoring provides a range of methods aimed at altering the syntactic structure of source code while preserving its underlying semantic information. Currently, code refactoring is being increasingly employed for code data augmentation since it can be utilized to automatically create additional source code samples. Therefore, we first chose code refactoring as one of the code data augmentation methods to generate the training candidate. Furthermore, we release all 18 atomic code refactoring operations adopted by GenCode.
Synonym Replacement (SR). The original SR is applied to text data. It augments a text by randomly selecting a certain number of words and then replacing the selected words with one of their randomly chosen synonyms. Here, synonyms are machine-generated. In contrast to text data, a program is composed of a series of statements. To augment the code data, we begin by randomly selecting a set number of statements from a program. Subsequently, each word within these selected statements is randomly replaced with one of its synonyms. All operations that may alter the syntactic structure follow the rule that preserves the original relationship between the features and the label in the code.
Random Insertion (RI). For text data, RI first randomly chooses a word in a sentence and then generates a number of synonyms of the chosen word. Lastly, RI randomly selects one of the generated synonyms and inserts it into a random position of this sentence. Generally, we repeat this operation $n$ times. In code data, we first randomly choose one of the statements and generate a number of synonyms for each word in this statement. Then, we randomly choose one word and insert one of its synonyms in a random position of the chosen statement.
Random Swap (RS). RS generally randomly chooses two different words in a sentence or two different sentences in a document and then swaps their positions. This operation is usually repeated $n$ times. For source code, we first randomly select two different statements in a program and then swap their positions, also we repeat it $n$ times.
Random Deletion (RD). RD randomly selects a certain number of words in a sentence or a certain number of sentences in a document and then deletes them. For source code, we randomly delete a certain number of words in a randomly chosen statement. Similarly, this method can preserve the original relationship between the features and the label in code data.
Back-translation (BT). BT translates the original text into another language and then translates it back to the original one to generate additional text data. For code data, similarly, we choose the English-French translation model Xie et al. (2020) bidirectionally for each statement in a program.
Not that: Different from Refactor and SR, atomic data augmentation operators including RI, RS, RD, and BT can break syntax and alter the label of the code. Despite this, we found that these code augmentation operators remain beneficial in GenCode for boosting the training of code models.
Appendix C Robustness Testing of Code Model
Robustness is an important property of DL models which indicates how the DL model can handle data with different distribution of the training data. Normally, there are two types of robustness of DL models, natural robustness and adversarial robustness. Natural robustness indicates how models perform when facing the natural distribution shift, for example, data in different environments and newly collected data. Adversarial robustness refers to the ability of models to defend the adversarial attacks. Robustness testing of code models has been addressed recently. Researchers proposed distribution shifted code (e.g., codes have the same functionality but are written by different programmers) data to study the natural robustness of code models. Besides, multiple adversarial attack methods have been proposed to test the adversarial robustness of code models, for example, Yang et al. (2022) proposed to change the variable name of code to force the model to make wrong predictions.
In this work, following existing works Yang et al. (2022), we employ adversarial attack methods proposed for code to evaluate the adversarial robustness of our trained model, namely, ALERT Yang et al. (2022) and MHM Zhang et al. (2020).
$ASR=\frac{\#Successfully\,generated\,adversarial\,data}{\#Corrected\,classified\,test\,data}\%$ (1)
In the evaluation of robustness, ASR (in Eq.( 1)) is the metric for assessing robustness. It calculates the percentage of successfully generated adversarial examples. Adversarial attacks are performed exclusively on correctly predicted test samples. A higher ASR suggests that an attack method is more effective, indicating lower robustness of the victim model.