### v2

# GenCode: A Generic Data Augmentation Framework for Boosting Deep Learning-Based Code Understanding

Abstract
Pre-trained code models lead the era of code intelligence with multiple models have been designed with impressive performance. However, one important problem, data augmentation for code data that automatically helps developers prepare training data lacks study in this field. In this paper, we introduce a generic data augmentation framework, GenCode, to enhance the training of code understanding models. Simply speaking, GenCode follows a generation-and-selection paradigm to prepare useful training code data. Specifically, it employs code transformation techniques to generate new code candidates first and then selects important ones as the training data by importance metrics. To evaluate the effectiveness of GenCode, we conduct experiments on four code understanding tasks (e.g., code clone detection) and three pre-trained code models (e.g., CodeT5). Compared to the state-of-the-art (SOTA) code augmentation method, MixCode, GenCode produces code models with 2.92% higher accuracy and 4.90% robustness on average.
Introduction
Deep learning (DL) for code has been one of the successful DL applications over the years. The methodology of applying deep learning to solve software engineering (SE) tasks, such as software vulnerability detection (Zhou et al. 2019), is commonly known as code learning. From the model level, different DL models specifically designed for code learning have been proposed, such as CodeBERT (Feng et al. 2020), GraphCodeBERT (Guo et al. 2020), and CodeT5 (Wang et al. 2021). From the commercial perspective, multiple related products such as Copilot  [^1] https://github.com/features/copilot and Tabnine  [^2] https://www.tabnine.com have been widely used to help our daily development. As a result, in different downstream learning-based code understanding tasks, e.g., code clone detection  (Lu et al. 2021) and program classification (Puri et al. 2021), DL already achieved significantly better results than traditional SE techniques.
However, most of the existing research focuses on designing new code representation techniques or code model architectures to further improve the effectiveness of DL for code, one important problem – how to prepare high-quality training data gained limited attention even though recent research demonstrated it highly affects the performance of code models (Sun et al. 2022). The challenge behind this is preparing training data for code models is heavy work. Although today developers can easily collect raw code data from the open-source community such as GitHub  [^3] https://github.com, data labeling is still expensive, especially for code-related tasks. It needs huge human effort with professional domain knowledge.
To tackle this problem, data augmentation is a famous technique to increase the diversity and quality of training data based only on existing labeled data. In the code learning field techniques such as code refactoring methods and text-oriented transformation methods can be directly used for code augmentation. Unfortunately, existing studies (Bielik and Vechev 2020; Yu, Wang, and Wang 2022) stated that code refactoring methods have limited advantages in the accuracy improvement of code models. Besides,  (Dong et al. 2023b) revealed that code models trained using existing data augmentation methods still suffer from the low robustness issue. Moreover, a recent work, MixCode (Dong et al. 2023a), that follows the spirit of Mixup (Zhang et al. 2018) to do code augmentation achieved promising results, but it can be only used for classification tasks. Therefore, there are currently no effective and reliable code augmentation methods that can be generalized across different tasks.
To bridge this gap, inspired by the great success of the typical two-stage Automated Data Augmentation (AutoDA) framework in the field of computer vision (Cubuk et al. 2020; Yang et al. 2023b), we propose GenCode, the first generation-and-selection-based data augmentation framework for code understanding. Specifically, at each training epoch, GenCode first generates training candidates by using different code transformation methods. Here, we consider different types of methods such as semantic-preserving code refactoring methods and syntax-breaking text transformation methods. After that, GenCode selects valuable training data from the candidates with the guidance of importance metrics. The important metric used in GenCode is simply defined as the loss value of code data produced by the code model. Finally, the selected samples are used for the model training. We conducted comprehensive experiments to study how GenCode can improve the accuracy and robustness of code models. The experimental results demonstrated that, compared to the SOTA method, MixCode, GenCode can produce code models with a 2.92% higher accuracy and 4.90% better robustness on average.
To summarize, the main contributions of this paper are:
  • We propose the first generation-and-selection-based framework, GenCode, which automatically enriches the training data for boosting the code understanding models.
  • We empirically demonstrate that GenCode can produce more accurate and robust code understanding models compared to existing data augmentation methods and conduct an ablation study to explore how different importance scores affect the effectiveness of GenCode.
Background and Related Work
Code Representation
Code representation is a crucial technique that converts source code into a format interpretable by Deep Neural Networks (DNNs), facilitating effective learning of source code features. Since source code can be seen as sequences of tokens or structural graph (Allamanis et al. 2018), researchers employ the techniques in Natural Language Processing (NLP) and Graph Neural Networks (GNNs) to learn information from code and solve downstream tasks, such as automated program classification (Puri et al. 2021) and code clone detection (Wang et al. 2020). Existing methods for code representation can be roughly divided into two groups, sequence-based representation and graph-based representation. Sequence-based representation follows the same procedure of text data processing in NLP to convert raw code data into a sequence of code tokens (Alon et al. 2019). Graph-based representation learns the structural information of code. Here, the structural information can be abstract syntax tree (AST), data flow information, and other types of flow information (Dinella et al. 2020; Yasunaga and Liang 2020).
Unlike code representation, our work focuses on code data preprocessing and studies the powerful pre-trained code models CodeBERT, GraphCodeBERT, and CodeT5 that achieved impressive results on code understanding tasks.
Data Augmentation for Code Understanding
Data augmentation is a well-studied technique in the machine learning field that increases the size and diversity of training data by only modifying the existing labeled data, thus, without further human effort. For instance, in the NLP field, researchers tried many ways to augment the text data, e.g., replacing a token with a synonym (Feng et al. 2021).
Recently, data augmentation has gained attention from software engineering researchers. Various data augmentation methods designed for source code have been proposed (Gao et al. 2020; Yefet, Alon, and Yahav 2020; Allamanis, Jackson-Flux, and Brockschmidt 2021; Pour et al. 2021; Bui, Yu, and Jiang 2021; Wang et al. 2022; Park, Kim, and Han 2023; Yang et al. 2023a), with the most general framework involving the use of code refactoring, e.g., local variable renaming and dead store, to create more code data without breaking the semantics of the original codes. Concretely, code refactoring rewrites the syntactic structure of source code while keeping the semantic information (Lacerda et al. 2020).
Existing works (Bielik and Vechev 2020; Yu, Wang, and Wang 2022; de Paula Rodrigues, Braga, and Dahab 2023; Dong et al. 2023b, 2024) conducted empirical studies and found code refactoring methods have limited advantages in boosting the training of code models. For instance, a recent empirical study (Dong et al. 2024) on code understanding revealed that the utilization of code refactoring methods only achieved success rates of 66.67% and 41.67% in improving accuracy and robustness, respectively, compared to code model training without data augmentation. Besides, the SOTA code data augmentation method, MixCode, which linearly interpolates pairs of source codes, has been proven more effective than the current general framework code refactoring. However, as a variant of Mixup specifically designed for code data, MixCode is limited to applications in code classification tasks, as it necessitates the label with one-hot encoding.
Compared to existing works, we propose a simple yet effective code augmentation framework that is not only unaffected by the dataset and model constraints but also applicable to both classification and non-classification tasks.
Definition
We first define the problem of data augmentation for code models.
Definition 1.
A program code  [^4] For simplicity, we use code in the paper. is defined as a 5-tuple $\mathcal{P}=\left(content,se,sy,Ip,Op\right)$: $content$ contains all the text in the code $\mathcal{P}$; $se$ and $sy$ are the semantic and syntax of the code $\mathcal{P}$; and $Ip$ and $Op$ are the input and output of the code $\mathcal{P}$.
Definition 2.
A code model is defined as a 3-tuple $M=\left(parameter,\mathcal{P},Y\right)$, where $parameter$ contains all the weights in code model $M$, $\mathcal{P}$ is the input code, and $Y$ is the label of $\mathcal{P}$ given a specific downstream code task.
Code models can be any machine learning model such as a simple regression model or a pre-trained model, e.g., CodeT5 (Wang et al. 2021).
Definition 3.
Given a code $P$, a code transformation method $\mathcal{T}:\mathcal{X}\rightarrow\mathcal{X^{\prime}}$, code augmentation is to transform $\mathcal{P}$ to $\mathcal{P^{\prime}}$ that $\mathcal{P^{\prime}}.content=\mathcal{T}\left(\mathcal{P}.content\right)\&\mathcal{P^{\prime}}.content\neq\mathcal{P}.content$.
Here, code transformation methods can be any method that can change the content of program statements such as text data transformation methods (e.g., synonym replacement and code refactoring methods (e.g., dead if statement adding).
Definition 4.
Given a code model $M$, a code $\mathcal{P}$ and its augmented version $\mathcal{P^{\prime}}$, semantic-preserving (breaking) code augmentation indicates $\mathcal{P}.se=\mathcal{P^{\prime}}.se$ $M(\mathcal{P})=M(\mathcal{P^{\prime}})$ ($\mathcal{P}.se\neq\mathcal{P^{\prime}}.se$ $M(\mathcal{P})\neq M(\mathcal{P^{\prime}})$).
Definition 5.
Given a code $\mathcal{P}$ and its augmented version $\mathcal{P^{\prime}}$, syntax-preserving (breaking) code augmentation indicates $\mathcal{P}.sy=\mathcal{P^{\prime}}.sy$ ($\mathcal{P}.sy\neq\mathcal{P^{\prime}}.sy$).
Assumption and Preliminary Study
GenCode selects the most valuable data for the code model training. This approach is grounded on the assumption that the data that code models have less confidence in is more useful for model training. To check whether this assumption is valid or not, we conducted a preliminary study to explore if there is a correlation between the accuracy of the trained code model and the confidence of the model on the training data. Specifically, the study is conducted on the Google Code Jam (GCJ) (Yang et al. 2022) dataset and CodeBERT pre-trained code model. The detailed steps are as follows:
Step 1. We initialize the code model and use code augmentation methods to generate additional data as the training candidate set. The code augmentation methods will be introduced in a later section.
Step 2. We randomly select a number of data from the training candidate set. Here the selected number is the same as the size of the original training data. We repeat this selection 100 times and collect 100 groups of training data. After that, we compute and record the average loss values for each group of training data.
Step 3. We train the code model using each group of training data respectively for a few epochs and thus save 100 trained code models. We then test each code model using the original test data and save 100 accuracies.
Step 4. We use the Pearson Correlation Coefficient (PCC) (Schober, Boer, and Schwarte 2018) metric to check the correlation between the loss values and the accuracies.
Table 1: Correlation between loss values and code model accuracy. The P-value less than 0.01 is highlighted with a blue background.
Coefficient P-value
Initialized (epoch 0) 0.6075 7.9687e-12
Middle (epoch 10) 0.2440 0.0125
Well-trained (epoch 20) 0.1657 0.0926
Table 1 summarizes the results of our preliminary study. Here, we considered three types of code models, the initialed model, the model that has been trained for a few epochs but not converged, and the well-trained code model, to check if our assumption can stand in different scenarios. The results indicate that there is a positive correlation between the loss values of data and the accuracy of code models trained by using these data. From the initialized model to the well-trained model, this correlation becomes weaker but still exists. We conjecture that when the model is in the early stage, most of the data that have high loss values are useful for the model training. However, when training proceeds, the data that still have high loss values are the model itself cannot learn. Thus, there will be a correlation drop from the early epoch to the later ones. Based on this phenomenon, we can propose an effective and reliable data augmentation framework to boost the training of code models based on this phenomenon.
Finding: There is a positive correlation between the loss values of code data and the accuracy of code models trained by using these data. When the code model is randomly initialized, this correlation is significant (with a P-value less than 0.01).
Methodology
Overview
The goal of GenCode is to produce code models with better performance on test data by training models on carefully generated data. Given a training set of program $\mathcal{P}_{train}$, its transformed version $\mathcal{P^{\prime}}_{train}$, and a set of test programs $\mathcal{P}_{test}$, after the training process $\mathsf{train}\left(\cdot\right)$ , $\varrho(M^{\prime}(\mathcal{P}_{test}))>\varrho(M(\mathcal{P}_{test}))$, where $M.parameter=\mathsf{train}(M,\mathcal{P}_{train})$ $M^{\prime}.parameter=\mathsf{train}(M^{\prime},\mathcal{P^{\prime}}_{train})$, and $\varrho\left(\cdot\right)$ is the performance measurement.
GenCode follows the generation-and-selection workflow and utilizes different types of code transformation methods to generate diverse code data first, then selects code data that are most useful for training as the training data. Algorithm 1 shows how GenCode works. In each training epoch, GenCode automatically generates more code data (Line 7 to Line 9, Definition 3) and then selects the most important ones (Line 10 to Line 13) to train the code model (Line 14). Here, the data importance is measured by the loss values produced by the code model which is the most common indicator to check how the model performs on the data. Overall, GenCode contains two main steps, utilizing code data augmentation methods to compose the search space and importance-guided data selection.
Algorithm 1 GenCode
1:$M$: initialized DNN model
2:$\mathcal{P},Y$: original code data and labels
3:$\mathcal{T}=\left\{t\right\}:$ a set of code augmentation methods
4:$M$: trained model
5:for $run\leftarrow\{0,\ldots,\#epochs\}$ do
6:     $\mathcal{P^{\prime}},Y^{\prime}\ \leftarrow\phi,\phi$triangleright initialization
7:     for $t\in\mathcal{T}$ do
8:         $\mathcal{P^{\prime}}\leftarrow\mathcal{P^{\prime}}\cup t\left(\mathcal{P}\right)$
9:         $Y^{\prime}\leftarrow Y^{\prime}\cup Y$triangleright code augmentation
10:     $I\leftarrow\mathsf{Importance}\left(M,\mathcal{P^{\prime}}\right)$triangleright computing importance
11:     $Index\leftarrow\mathsf{argSort}\left(I\right)\left[:||\mathcal{P}||\right]$triangleright sort by importance
12:     $X_{train}\leftarrow\mathcal{P^{\prime}}\left[Indext\right]$
13:     $Y_{train}\leftarrow Y^{\prime}\left[Indext\right]$
14:     $M\leftarrow\mathsf{Train}\left(M,X_{train},Y_{train}\right)$
15:return $M$
Search Space
Even though the code has strict syntax constraints, recent work (Dong et al. 2023b) found that these data augmentation methods that slightly break the syntax of the source code e.g., Random Swap, can still be useful for code model training. To enhance the diversity of source code within the feature space, therefore, GenCode considers two types of code augmentation methods, semantic/syntax-preserving methods, and syntax-breaking methods (see Definition 4 and Definition 5).
GenCode adapts code refactoring methods as semantic/syntax-preserving methods. Code refactoring methods follow the constraints of programming languages to edit the code while keeping the main functionality of the code. On the other hand, text augmentation methods are used as syntax-breaking methods in GenCode which are widely used in the NLP field. For example, text augmentation methods can randomly delete some tokens in the code. However, even though the syntax has been broken, the code model can still successfully predict those codes since it reads the code as a sequence. In total, as shown in Table 2, GenCode currently supports 18 code refactoring methods and 5 text augmentation methods. Besides, Figure 1 provides an example of the training candidate that is to generate the search space. In the syntax-preserving, we present a code refactoring method named API Renaming, which involves modifying the function name from ”main()” to ”even()”. In the syntax-breaking, we demonstrate another instance of API Renaming, where two different statements within a program are randomly selected and their positions swapped. As depicted in this example, the third and fourth lines have been interchanged.
Table 2: Atomic data augmentation operators adopted in GenCode.
Syntax-preserving Syntax-breaking
API renaming
Filed enhancement
Synonym replacement (SR) Random insertion (RI) Random swap (RS Random deletion (RD) Back-translation (BT)
Arguments adding
For loop enhancement
Arguments renaming
If enhancement
Dead for adding
Local variable adding
Dead if adding
Local variable renaming
Dead if-else adding
Method name renaming
Dead switch adding
Plus zero
Dead while adding
Print adding
Duplication Return optimal
Figure 1: An example of atomic data augmentation operators.
Table 3: Effectiveness of data augmentation methods w.r.t. test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on original test data and ASR $\downarrow$ (%) on test data. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
GCJ Refactory Java250 BigCloneBench
Model DA method Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$
No Aug 90.96 ± 0.07 33.31 56.13 96.21 ± 0.13 36.31 42.31 96.36 ± 0.09 40.55 53.37 96.91 ± 0.12 9.74 24.61
Refactor 91.74 ± 0.11 35.56 52.84 95.95 ± 0.19 32.64 40.89 96.44 ± 0.12 38.35 52.56 96.97 ± 0.21 8.55 18.29
SR 70.66 ± 0.11 35.46 52.71 96.71 ± 0.11 39.25 53.91 96.21 ± 0.02 42.01 53.79 97.02 ± 0.15 12.52 26.11
RI 59.91 ± 0.21 40.78 65.76 96.43 ± 0.18 38.45 45.57 96.15 ± 0.06 43.15 55.07 97.38 ± 0.23 8.31 30.08
RS 93.24 ± 0.13 28.26 52.44 96.73 ± 0.15 38.63 45.98 96.49 ± 0.04 39.36 50.79 96.92 ± 0.12 13.41 30.13
RD 76.98 ± 0.09 35.65 55.46 96.46 ± 0.18 30.23 41.32 96.61 ± 0.12 38.31 52.42 97.11 ± 0.19 9.58 30.15
BT 82.74 ± 0.15 40.38 55.95 94.31 ± 0.38 62.98 69.87 96.06 ± 0.06 42.21 53.32 97.13 ± 0.17 11.29 26.32
MixCode 92.21 ± 0.25 29.28 52.53 96.76 ± 0.31 31.04 41.01 96.62 ± 0.23 36.14 50.67 - - -
CodeBERT GenCode 96.23 ± 0.18 22.33 46.62 99.32 ± 0.14 23.75 31.25 98.89 ± 0.09 29.89 40.65 98.93 ± 0.11 7.12 15.83
No Aug 93.96 ± 0.15 31.23 56.02 96.84 ± 0.16 22.97 30.26 96.41 ± 0.11 23.16 42.66 96.88 ± 0.09 4.19 12.22
Refactor 91.72 ± 0.21 30.31 56.58 95.49 ± 0.21 24.26 32.57 96.53 ± 0.15 23.13 42.42 97.11 ± 0.23 6.05 6.17
SR 89.46 ± 0.13 40.02 66.68 95.57 ± 0.13 28.39 47.15 96.46 ± 0.14 22.92 42.18 97.14 ± 0.16 6.13 10.16
RI 80.47 ± 0.21 43.51 65.75 94.51 ± 0.25 25.08 32.87 96.52 ± 0.09 22.37 42.33 97.35 ± 0.17 2.23 6.28
RS 94.76 ± 0.15 30.18 54.75 97.88 ± 0.15 21.85 29.67 96.44 ± 0.14 22.68 41.98 97.54 ± 0.12 6.03 11.94
RD 91.71 ± 0.26 31.73 60.92 96.53 ± 0.19 20.35 26.98 96.56 ± 0.21 22.39 41.84 97.16 ± 0.21 6.18 8.27
BT 80.47 ± 0.05 41.65 62.99 95.99 ± 0.27 35.56 46.23 96.39 ± 0.13 22.91 42.32 97.11 ± 0.16 8.18 14.12
MixCode 94.19 ± 0.25 30.26 54.49 97.29 ± 0.15 14.89 26.53 96.89 ± 0.19 20.03 42.01 - - -
GraphCodeBERT GenCode 97.93 ± 0.14 26.43 48.87 99.45 ± 0.17 15.78 21.92 99.16 ± 0.13 18.76 33.81 99.23 ± 0.08 1.86 4.62
No Aug 95.02 ± 0.07 30.78 55.67 97.22 ± 0.17 22.45 29.69 96.98 ± 0.08 31.76 49.82 97.38 ± 0.13 10.32 22.97
Refactor 92.21 ± 0.14 30.12 55.93 96.43 ± 0.22 28.76 35.45 96.92 ± 0.11 30.66 47.12 97.24 ± 0.12 9.32 22.05
SR 90.14 ± 0.21 39.21 51.84 94.67 ± 0.14 28.24 36.26 96.53 ± 0.05 28.65 46.01 97.31 ± 0.08 15.87 25.98
RI 78.77 ± 0.05 41.02 63.46 95.02 ± 0.15 24.78 33.61 96.11 ± 0.09 27.14 46.76 97.46 ± 0.13 10.43 29.76
RS 95.66 ± 0.14 29.35 51.74 97.97 ± 0.29 21.32 27.76 96.34 ± 0.11 28.45 48.45 97.03 ± 0.15 14.33 30.43
RD 92.01 ± 0.11 31.04 56.85 97.52 ± 0.26 19.93 24.87 97.21 ± 0.07 26.09 45.24 97.43 ± 0.07 9.98 28.87
BT 81.45 ± 0.13 39.65 54.24 95.38 ± 0.11 33.62 43.67 96.45 ± 0.14 29.47 45.38 97.65 ± 0.14 10.43 25.34
MixCode 94.32 ± 0.16 29.82 42.56 97.32 ± 0.13 12.49 18.32 96.91 ± 0.11 28.73 48.87 - - -
CodeT5 GenCode 98.85 ± 0.11 23.57 44.76 99.64 ± 0.06 12.13 19.56 99.38 ± 0.11 21.63 39.76 99.34 ± 0.12 7.39 12.15
Importance-Guided Data Selection
After constructing the training candidate, we then choose the most valuable code data that can benefit the code model training. Based on our Finding drawn from the preliminary study, GenCode chooses the simplest way to measure the importance of code data to the model – the loss value. That means, GenCode takes all the transformed code as input to the code model and records all the loss values first, then ranks all the code in descending order based on the loss values and selects the first $K$ code as the training data. Following the common setting in the data augmentation problem (Dong et al. 2023a), $K$ is set as the number of original training data. GenCode chooses loss value as the indicator since it can be generalized to any code model and task.
Evaluation
We evaluate both the accuracy and robustness of trained code models. Besides, we check how does importance score affects the effectiveness of GenCode.
Datasets and Models
Datasets and Models We evaluate GenCode on four code understanding tasks, bug detection (Hu et al. 2019), authorship attribution (Yang et al. 2022), code clone detection (Yang et al. 2022) (Note that, MixCode is not suitable for this task since its input is code pairs.), and problem classification (Puri et al. 2021). For each code task, we fine-tune three types of models based on pre-trained code models, CodeBERT, GraphCodebERT, and CodeT5. For the model training, we follow existing works (Hu et al. 2019; Wang et al. 2021; Puri et al. 2021; Yang et al. 2022; Dong et al. 2023a), to set all configurations. Specifically, we set the training epoch as 50 for all tasks and models. Regarding the optimizer, we utilize Adam (Kingma and Ba 2014) with a learning rate of $10^{-3}$ for all the models. Besides, we use early stopping with a patience of 20 to mitigate overfitting. To reduce the influence of randomness, we train each model five times and report the average results along with the standard deviation. For MixCode, we follow the recommendation of the original paper and set the Mixup ratio as $\alpha=0.1$. For the setting of adversarial attacks, we directly use the default configurations provided by the original papers  (Yang et al. 2022).
Evaluation Metrics. In the evaluation, we assess the performance of trained models from two perspectives, namely, clean accuracy and model robustness. For the robustness evaluation, we check the attack success rate (ASR) metric of two well-established adversarial attack methods, namely, ALERT (Yang et al. 2022), which uses variable renaming while considering natural semantics, and MHM (Zhang et al. 2020), which also randomly replaces variable names but does not account for natural semantics. We do not consider the natural robustness since the existing natural robustness datasets can not fully support our studied tasks, i.e., no natural robustness datasets support authorship attribution, bug detection, and code clone detection.
Baseline Methods. We consider model training without data augmentation (No Aug), Refactor (Yu, Wang, and Wang 2022), a family of text-oriented data augmentation methods (Dong et al. 2023b), and MixCode as baseline methods. Refactor, which keeps code syntax rules, is a widely used data augmentation framework specifically designed for code data. In contrast, text-oriented code data augmentation methods, which slightly break syntax rules, have recently proven more effective than Refactor. Additionally, MixCode is a more recent and high-performing code data augmentation method that significantly enhances code models.
Accuracy Evaluation
Overall accuracy. Column Accuracy in Table 3 summarizes the final test accuracy of code models trained by using different code data augmentation methods including GenCode. Surprisingly, from the results, we can see that most of the existing code augmentation methods except RS and MixCode negatively affect the performance of code models. This finding is similar to the conclusion from recent code augmentation studies (Bielik and Vechev 2020; Yu, Wang, and Wang 2022; Dong et al. 2023b) which states code augmentation has limited advantage in boosting the training of code models and also indicates that data augmentation for code is a challenging problem.
Then, compared to the existing data augmentation methods, the results demonstrate that GenCode consistently enhances the training and achieves the best results across all the considered datasets and models. Concretely, on average, GenCode generates models with 4.52%, 3.05%, 2.56%, and 2.11% higher accuracy than No Aug in GCJ, Refactory, Java250, and BigCloneBench, respectively. Besides, compared to MixCode, GenCode can achieve 4.1%, 2.34%, and 2.33% better results in GCJ, Refactory, and Java250 datasets, respectively.
Statistical analysis. We conducted statistical tests, including the T-Test and Wilcoxon signed-rank test, across all pre-trained models. The results show that GenCode significantly (with p-value $<$ 0.05) outperforms other methods including MixCode. In conclusion, GenCode is effective in training accurate code models.
(a) Authorship attribution
(b) Bug detection
(c) Problem classification
(d) Clone detection
Figure 2: Convergence speed of CodeBERT using different code data augmentation methods in each task.
Convergence speed analysis. Figure 2 depicts the average test accuracy at each training epoch of CodeBERT using different data augmentation methods. The results demonstrated that GenCode consistently outperforms other baseline methods during the whole training process. We can conclude that the choice of the number of training epochs will not affect the advantages of our method over other methods.
(a) Refactor
(b) SR
(c) RI
(d) RS
(e) RD
(f) BT
(g) MixCode
(h) GenCode
Figure 3: Visualization of code embeddings after dimension reduction using Principal Component Analysis (PCA). Model: CodeBERT, dataset: Refactory, task: Bug detection.
Considering Figure 3, we can see that the embedding produced by GenCode is clearly distributed compared to other methods. Besides, we computed the average distance between the data points from different classes, i.e., the average distance between red points and blue points in Figure 3. The results are shown in the following Table. The results confirm that GenCode can better distinguish the two classes. Interestingly, we found that MixCode has the smallest distance result. That means even though MixCode can train code models with good performance, the code models have low confidence in their predictions, thus, the data are near the decision boundaries. This phenomenon is similar to the findings in (Verma et al. 2019), where the representations obtained by interpolating hidden states have two desirable properties: the class-representations are flattened into a minimal amount of directions of variation, and all points in between these flat representations, most unobserved during training and off the data manifold, are assigned low-confidence predictions.
Table 4: Average distance between the data points from different classes. The best is highlighted with a blue background, and the worst is highlighted with a gray background.
DA method Refactor SR RI RS
Average distance 6.02 8.05 7.02 8.06
DA method RD BT MixCode GenCode
Average distance 7.03 6.00 4.05 8.78
Table 5: Mean values of maximum probabilities. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
Model DA method GCJ Refactory Java250 BigCloneBench
No Aug 0.0381945 0.9368535 0.9554731 0.9651272
Refactor 0.0391864 0.9643155 0.9620814 0.9654681
SR 0.0380562 0.9527686 0.9341456 0.9583812
RI 0.0371215 0.9189242 0.9327171 0.9704512
RS 0.0392987 0.9561591 0.9581767 0.9410937
RD 0.0392987 0.9723022 0.9634647 0.9631375
BT 0.0379581 0.9566747 0.9417861 0.9667154
MixCode 0.0390714 0.9687351 0.9685594 -
CodeBERT GenCode 0.0413779 0.9808122 0.9714346 0.9786421
Table 6: Effectiveness of data augmentation methods w.r.t. test accuracy $\uparrow$ (average $\pm$ standard deviation, %) on original test data and ASR $\downarrow$ (%) on test data. The best results are highlighted in blue. Code understanding tasks include Authorship attribution (GCJ), Bug detection (Refactory), Problem classification (Java250), and Clone detection (BigCloneBench).
GCJ Refactory Java250 BigCloneBench
Model DA method Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$ Accuracy $\uparrow$ MHM $\downarrow$ ALERT $\downarrow$
GenCode (Random) 94.74 ± 0.21 25.21 50.32 98.82 ± 0.15 27.65 36.65 98.24 ± 0.11 32.52 44.87 98.27 ± 0.19 8.01 18.12
GenCode (Min) 93.34 ± 0.32 35.66 66.67 96.13 ± 0.21 65.51 69.76 96.43 ± 0.17 46.76 53.86 97.23 ± 0.14 20.65 41.32
CodeBERT GenCode 96.23 ± 0.18 22.33 42.62 99.32 ± 0.24 22.75 29.25 98.89 ± 0.09 29.89 40.65 98.93 ± 0.15 7.12 15.83
GenCode (Random) 95.26 ± 0.11 29.97 50.63 98.93 ± 0.18 17.56 24.29 98.68 ± 0.22 20.65 38.86 98.67 ± 0.13 2.04 5.97
GenCode (Min) 93.53 ± 0.17 40.55 68.23 96.31 ± 0.24 38.73 46.98 96.72 ± 0.12 26.87 46.67 97.22 ± 0.21 11.45 21.23
GraphCodeBERT GenCode 97.93 ± 0.14 26.43 39.87 99.45 ± 0.17 15.78 20.92 99.16 ± 0.13 18.76 33.81 99.23 ± 0.09 1.86 4.62
GenCode (Random) 96.24 ± 0.21 25.03 47.61 98.99 ± 0.21 18.46 23.67 98.93 ± 0.16 25.11 42.34 98.83 ± 0.08 8.67 17.42
GenCode (Min) 94.11 ± 0.31 38.65 59.89 98.29 ± 0.12 35.67 48.59 97.03 ± 0.23 32.61 49.27 97.01 ± 0.15 18.36 25.54
CodeT5 GenCode 98.85 ± 0.16 23.57 38.76 99.64 ± 0.37 12.13 18.81 99.38 ± 0.11 21.63 39.76 99.34 ± 0.12 7.39 12.15
Visualizing principal component analysis. Furthermore, we explore why models trained by GenCode have higher accuracy than models trained using other methods by analyzing the code representation capability of trained code embeddings. Specifically, 1) first, we extract the code embeddings produced by each trained model using all the test data as code features. 2) Then, since the extracted features are high dimensions and hard to analyze, we reduce the dimensions of code data features using the Principal Component Analysis (PCA) (Maćkiewicz and Ratajczak 1993) algorithm. 3) After that, we visualize the reduced 2-dimensional vectors. Normally, the features extracted by a well-trained model are more clearly distributed based on the data labels. Figure 3 depicts the visualized features extracted from CodeBERT using the Refactory dataset. We can see that the code embeddings trained by using GenCode are more clearly distributed than embedding extracted by other models, i.e., the buggy code and correct code are greatly different in the figure. These results indicate that GenCode can train good code embeddings and, therefore, produce better accurate code models for code understanding tasks. Besides, we computed the average distance between the data points from different classes, i.e., the average distance between red points and blue points in Figure. 3. The results from Table 4 confirm that GenCode can better distinguish the two classes. Interestingly, we found that MixCode has the smallest distance result. That means even though MixCode can train code models with good performance, the code models have low confidence in their predictions, thus, the data are near the decision boundaries. This phenomenon is similar to the findings in (Verma et al. 2019), where interpolating hidden state representations yields two beneficial properties: 1) class representations are flattened into a minimal number of directions of variation, and 2) all points between these flat representations are assigned low-confidence predictions.
Robustness Evaluation
Overall robustness. Column MHM and ALERT in Table 3 summarize the attack success rate of adversarial attack methods against code models. First, we can see that similar to the observation in Accuracy Evaluation, not all data augmentation methods can improve the robustness of code models. This finding is consistent with the existing works (Dong et al. 2023b). Specifically, in the baseline methods, only RS and MixCode have a positive impact on the robustness of code models. Considering our method, GenCode consistently improves the robustness of code models and achieves the best results in 21 out of 24 cases. On average, compared to No Aug, models trained by GenCode have an 8.42% lower ASR (higher robustness). This result breaks the previous finding (Bielik and Vechev 2020) that states simply adding additional generated code to the training dataset to train code models has limited benefits in improving their robustness. On the other hand, compared to MixCode, GenCode can train code models with 4.90% higher robustness on average.
Statistical analysis. We also conduct statistical tests using T-Test and Wilcoxon signed-rank test. When comparing existing data augmentation methods, GenCode consistently exhibits a p-value $<$ 0.05 across all pre-trained models. In conclusion, GenCode can produce more adversarial robust code understanding models than other methods.
Maximum probability analysis. To explore why GenCode produces a more robust code model, we check the confidence of trained models using different code data augmentation methods on the original test data. Normally, if the DNN model is more confident in the test data, it is more difficult to attack this model using this data (Wu et al. 2018). To do so, for each dataset and code model, we collect the maximum probability of the corrected classified test data produced by the code model first. Then, we calculate the mean values of these maximum probability scores. The results of the mean values of maximum probabilities are shown in Table 5. We can see that code models trained by GenCode are more confident in the test data than models trained using other code data augmentation methods. As a result, GenCode trained code models are harder to attack.
Ablation Study
Even though we have preliminarily shown that selecting the code samples with higher loss values is a valid way to train good models, it is still necessary to study to what extent the advantage can be brought from this selection strategy. Therefore, in this part, we conduct an ablation study to compare GenCode with its two variants, GenCode with data selection using minimum loss values - GenCode (Min) and random selection - GenCode (Random), respectively.
Table 6 summarizes the accuracy and robustness of trained code models using different selection strategies. We can see GenCode consistently outperforms the two variants under all testing criteria. Selecting code samples with minimum loss values is the worst way for GenCode. Interestingly, we found that GenCode (Random) performs better than all baseline methods (i.e., No Aug, Refactor, and MixCode) in terms of both accuracy and robustness improvements based on the results. This phenomenon indicates simply combining all syntax-preserving methods including code refactoring methods and syntax-breaking including text-oriented methods is already a good code data augmentation method.
Discussion
Even though the accuracy of our studied code models already achieved relatively high accuracy in these tasks, i.e., most of them have more than 90% accuracy, their robustness is still low. Regarding the large language models (LLMs), we have checked the literature and found that the clone detection task is still challenging for LLMs, e.g., ChatGPT has only 0.784 Precision for Java-Java pair detection (Khajezade et al. 2024). For other program understanding tasks that have a large number of classes, e.g., the Java250 dataset used in our experiments has 250 classes. It is also hard to use open-source LLMs such as ChatGPT to predict the label since we need to add many label guidance in the prompt.
Conclusion
We propose the first generation-and-selection-based framework, GenCode, to boost code understanding models. The basic assumption behind GenCode is that the less confident data contributes more to the model training. Based on this assumption, GenCode utilizes atomic data augmentation operators to generate candidate training data first and selects data with maximum loss values to train the code model. Our evaluation demonstrates that GenCode can produce more accurate and robust models than MixCode with 2.92% accuracy and 4.90% robustness improvement, respectively. Besides, we analyze how the importance score affects the effectiveness of GenCode.
In the future, we plan to 1) explore the combination of GenCode with other confidence scores, such as uncertainty metrics and active learning methods; 2) study the effectiveness of GenCode on fine-tuning open-source large language models, e.g., Llama.